OpenStudy (anonymous):

Max tossed a fair coin 3 times. What is the probability of getting heads in the first two trials and tails in the last trial? A. 4/8 B. 3/8 C.2/8 D.1/8

OpenStudy (ronlover101):

it is either b or c

OpenStudy (ronlover101):

but i am leaning towards b

OpenStudy (gabebae):

He's doing three trials

OpenStudy (gabebae):

I'd go with B as well

OpenStudy (ronlover101):

I am meaning a or b

OpenStudy (ronlover101):

me two

OpenStudy (gabebae):

@nmcclure17 Pick B

OpenStudy (anonymous):

okay thanks What are you going to say @ospreytriple ?

OpenStudy (anonymous):

Since each toss is an independent event, you multiply the probabilities for each together, i.e.\[P(HHT)=P(H)\times P(H)\times P(T)\] And you probably know the probability of getting a head or a tail on one coin flip.

OpenStudy (anonymous):

ya its 1/2

OpenStudy (anonymous):

Nope. Try again. What's the probability of getting a head on the first toss?

OpenStudy (anonymous):

1/3?

OpenStudy (anonymous):

Sorry, I might have misunderstood you last comment. You were meaning that the probability of getting a head (or the probability of getting a tail) on each toss is 1/2?

OpenStudy (anonymous):

If that's what you meant, then you're correct. Now multiply the three probabilities together to get the overall probability.

OpenStudy (anonymous):

1/8

OpenStudy (anonymous):

@ospreytriple ??

OpenStudy (anonymous):

You're correct. Good job. Sorry, I lost my connection.

OpenStudy (anonymous):

oh its okay thanks :)